OpenStudy (anonymous):

Which one of these equations is correctly written in Standard Form? one over fourx + y = −6 4x + 3.5y = −6 x + 4y = −24 −4x + y = −24

OpenStudy (anonymous):

standard from is Ax +By = C 1 -- x + y = - 6 4

OpenStudy (anonymous):

just cancel out the fraction

OpenStudy (anonymous):

thank you so much heres a medal

OpenStudy (anonymous):

to do that an EXAMPLE of canceling out a fraction i dont want 2 there 1 2 [ -- x + y = 8 ] multiply whole equation by 2 2 what u do to 1 side u do to the other 1 x + 2 = 16 standard form

OpenStudy (anonymous):

thanks